OpenStudy (anonymous):

Provide the complete solution for 3x-2 = 7 (mod 11) 5 = 4x-1 (mod 9) Provide the complete solution for 3x-2 = 7 (mod 11) 5 = 4x-1 (mod 9) @Mathematics

OpenStudy (lilg132):

5 + 1 = 4x 6 = 4x 1.5 = x (mod 9 answer) 3x - 2 = 7 3x = 7 + 2 3x = 9 x = 3

OpenStudy (lilg132):

(mod 11 was the second answer)

OpenStudy (lilg132):

remember to click good answer :D

OpenStudy (anonymous):

I don't understand your answer, is this the complete solution to the linear congruences?